Problem 12-3 Return Calculations [LO1] Suppose a stock had an initial price of $60 per share, paid a dividend of $2.20 per share during the year, and had an ending share price of $49. Compute the percentage total return. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Total return % What was the dividend yield and the capital gains yield? (Negative amounts should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) Dividend yield % Capital gains yield %
Answer #1

Total return=(End price-Beginning price+Dividends)/Beginning price

=(49-60+2.2)/60=(14.67%)(Approx)(Negative).

Dividend yield=Dividend/Beginning price

=(2.2/60)=3.67%(Approx).

Capital gains yield=(End value-Beginning price)/Beginning price

=(49-60)/60=(18.33%)(Approx)(Negative).